OpenStudy (anonymous):

Using addition elimination method please solve this: 6c - 9d = 103 and 5c - 9d = 103 please show work.

OpenStudy (anonymous):

Sorry the first equation equals 111.

OpenStudy (mathmate):

6c - 9d = 111 5c - 9d = 103 Subtract second from first to get c=8 Substitute c=8 in equation 1 to get 9d=48-111=-63 or d=-7 Substitute both c and d in second equation to check answer.

jhonyy9 (jhonyy9):

- ok so for using addition elimination method so for this you need firstly multiply the second equation by -1 and hence will get for second equation -5c+9d=-103 - so now adding these equations will get 6c-9d=111 -5c+9d= -103 -------------- (6c+(-5c)+(-9d+9d)=111+(-103) (6c-5c)+0=111-103 c+0=8 c=8 so than 6*8-9d=111 so 48-9d=111 so -9d= 111-48 so -9d=63 so d= 63/(-9) so d= -63/9 = -7 than we get c=8 d= -7